Find an equation of the line through the pair of points. Write the equation in the form Ax + By = C. (8,-6) and (4, -3)
Answer by stanbon(75887)

Find an equation of the line through the pair of points. Write the equation in the form Ax + By = C. (8,-6) and (4, -3)
-----------------
slope = (-3--6)/(4-8) = 3/-4 = -3/4
----
Using form y = mx +b, solve for "b":
-3 = (-3/4)(4) + b
b = 0
----------------------------
Equation:
y = (-3/4)x
-3x - 4y = 0
OR
3x + 4y = 0
=================
Cheers,
Stan H.